OpenStudy (anonymous):

The incomplete work of a student to solve an equation is shown below: Step 1: 4x + 12 = 4 Step 2: ? Step 3: x = −8 ÷ 4 Step 4: x = −2 What is the missing Step 2? 4x = 8 4x = 16 4x = −16 4x = −8

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (mathmale):

How would YOU solve this equation for x? Try it, please.

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (study312):

Well if part 3 is x = -8 ÷ 4 then part 2 is 4x = -8. It would be 4x + 12 -12 = 4 - 12, = 4x = -8. I pretty sure this is right

OpenStudy (study312):

Get it?

OpenStudy (anonymous):

so yea its D

OpenStudy (study312):

Should be.